Question 5 1 pts A tailor must make 6 of the same size suits for a client. The first suit took 13.2 hours to make and the last suit took 7.6 hours to make what was the tailors learning curve % of this run of suits assuming a steady state of production is never reached? Express your answer in 96 to the nearest 0.1%
Answer #1

Let rate of learning curve be r. This means the tailor takes r% less time each year.

7.6 = 13.2 (1 - r/100)6

7.6/13.2 = (1 - r/100)6

1 - r/100 = 0.91

r/100 = 0.09

r = 9%

The answer is 9.0%.